Determine the solution:
5/4x-2/3y=3; 1/4x+5/3y=6

Answer by eshomali(33)

Possible Solutions for X and Y:
1. X=(15/8)Y + 24 ; Y=(-16/30)X + 4.5
2. X=(-20/3)Y + 24 ; Y=(-3/20)X + (18/5)
:D